An investor is considering the purchase of Gryphon stock, which has returns given in the table below.

Scenario

Probability

Rate of Return

Recession

0.2

-2%

Normal Economy

0.54

5%

Boom

0.26

13%

Calculate the expected return and standard deviation of Gryphon. Round your answers to 2 decimal places.

Enter your answers below.

E(r) =     %

Std. Dev. =   %
Answer #1

Expected return=Respective return*Respective probability

=(0.2*-2)+(0.54*5)+(0.26*13)=5.68%

probability Return probability*(Return-Expected return)^2
0.2 -2 0.2*(-2-5.68)^2=11.79648
0.54 5 0.54*(5-5.68)^2=0.249696
0.26 13 0.26*(13-5.68)^2=13.931424
Total=25.9776%

Standard deviation=[Total probability*(Return-Expected return)^2/Total probability]^(1/2)

=(25.9776)^(1/2)

=5.1%(Approx).